OAU Post-UTME Mathematics 2021 (Batch B) Question 2

Question 2

Probability

There are 6 blue and 4 red identical balls in a bag. Two balls are drawn at random one after the other without replacement. What is the probability that the two balls are of the same colour?

  1. A.26/45
  2. B.7/15
  3. C.12/25
  4. D.½
Show the answer and explanation

Answer: B (7/15)

Explanation

P(both red) = 4/10 × 3/9 = 12/90 and P(both blue) = 6/10 × 5/9 = 30/90. Add: 42/90 = 7/15.